Compute the median, mode, and mean for 1,2,3,3,3,4,5 with the data toolpak by choosing Descriptive Statistics. Are the answers the same as what you did manually?
Answer #1

Using the Data Tool pack the Mean ,Median , Mode can be calculated as:

Now,Manually Mean,Median,Mode can be calculated as:

Mean=\frac{\sum x_i}{n}=\frac{21}{7}=3

Median can be calculated as:

Since our n is odd Thus, Median is:

Median=[\frac{n+1}{2}]^{th} obs=[\frac{7+1}{2}]^{th }obs.=4^{th} obs=3

Now,Mode can be caculated as:

Observation 3 has the highest frequency,Thus,

Mode=3

Thus,It is conclude that Answer are the same.

If you have any difficulty to understand the problem please comment. Please rate the answer.